OpenStudy (anonymous):

Someone help i have no idea how to do this Which shows 54^2 - 46^2 being evaluated using the difference of squares method? 54^2 - 46^2 = (2916 + 2116)(2916 - 2116) = 4,025,600 54^2 - 46^2 = (54 + 46)(54 - 46) = (100)(8) = 800 54^2 - 46^2 = 2916 - 2116 = 800 54^2 - 46^2 = (54 - 46)^2 = 8^2 = 64

OpenStudy (anonymous):

i would go with this one \[54^2 - 46^2 = (54 + 46)(54 - 46) = (100)(8) = 800\]

OpenStudy (anonymous):

\[a^2-b^2=(a+b)(a-b)\] is what is being used

OpenStudy (anonymous):

thank you so much ^_^